### v1

# Evaluating Knowledge Graph Based Retrieval Augmented Generation Methods under Knowledge Incompleteness

Abstract
Knowledge Graph based Retrieval-Augmented Generation (KG-RAG) is a technique that enhances Large Language Model (LLM) inference in tasks like Question Answering (QA) by retrieving relevant information from knowledge graphs (KGs). However, real-world KGs are often incomplete, meaning that essential information for answering questions may be missing. Existing benchmarks do not adequately capture the impact of KG incompleteness on KG-RAG performance. In this paper, we systematically evaluate KG-RAG methods under incomplete KGs by removing triples using different methods and analyzing the resulting effects. We demonstrate that KG-RAG methods are sensitive to KG incompleteness, highlighting the need for more robust approaches in realistic settings.
1 Introduction
Large Language Models (LLMs) have achieved remarkable success across various natural language processing tasks such as Question Answering (QA) (Achiam et al., 2023; Touvron et al., 2023; Jiang et al., 2023). However, they face several limitations, including outdated knowledge (Dhingra et al., 2022), insufficient domain-specific expertise (Li et al., 2023), and a tendency to generate plausible but incorrect information, known as hallucination (Ji et al., 2023).
Retrieval-Augmented Generation (RAG) mitigates these issues by integrating information retrieval mechanisms, allowing LLMs to access up-to-date and reliable information without requiring modifications to their architecture or parameters (Khandelwal et al., 2019; Ram et al., 2023; Borgeaud et al., 2022; Izacard & Grave, 2021). While conventional RAG approaches rely on unstructured text corpora, recent studies (Chen et al., 2023; Han et al., 2024; Zhang et al., 2025) have started to explore knowledge graphs (KGs) as knowledge resources by exploring different retrieval methods, such as neighborhood extraction after entity linking and similarity-based entity search in the embedding space, as well as different prompting methods like knowledge re-writting (Wu et al., 2024). Such KG-based RAG (KG-RAG for shortness) methods can reduce text redundancy (Li et al., 2024), enable flexible updates (Paulheim, 2016), and provide structured reasoning evidence (Luo et al., 2024). Despite these benefits, KGs are often incomplete in real-world applications (Min et al., 2013; Ren et al., 2020; Pflueger et al., 2022), raising a crucial question: Can KG-RAG methods remain effective when the given KG is incomplete?
LLMs have demonstrated substantial knowledge retention and reasoning capabilities Achiam et al. (2023); Guo et al. (2025). Consequently, KG-RAG methods (Luo et al., 2024; Sun et al., 2024; He et al., 2024) are expected to leverage LLMs to infer relevant facts that extend beyond explicitly retrieved triples. For instance, from the retrieved triples and , LLMs can infer that Jaxon Bieber is Justin Bieber’s brother by recognizing the underlying semantic path. In this way, KG-RAG methods can effectively compensate for missing direct evidence and mitigate the impact of incomplete KGs.
However, existing KG-RAG studies directly adopt KG QA benchmarks (Yih et al., 2016; Talmor & Berant, 2018) for evaluation. In these benchmarks, the ground truth answer of each question can be directly inferred from the knowledge in the KG, failing to reflect real-world scenarios where KGs are incomplete for answering the question. Thus the current evaluation does not reflect KG-RAG methods’ performance facing incomplete knowledge. In this paper, we systematically assess three popular KG-RAG methods under varying levels of KG incompleteness. Our findings reveal that these methods still benefit significantly from incomplete KGs, but are sensitive to missing knowledge with performance drops, emphasizing the need for more robust approaches for KG-RAG.
2 Methodology
2.1 Background and Notations
Knowledge Graph.
Given two finite sets and , whose elements are called entities and relations, a KG is a directed multi-relational graph represented as a subset of , where each element is called a triple.
Reasoning Path.
A reasoning path in a KG is a directed sequence of entities and relations that connect a source entity to a target entity, forming a logical pathway for inference:
where , , and denotes the length of the reasoning path. For example, given the question “Who is the brother of Justin Bieber?”, a possible reasoning path to the answer, as shown in Figure 1, is:
2.2 Benchmark Datasets
We build our experimental datasets based on two KGQA benchmarks that are mostly used by current KG-RAG methods: WebQuestionsSP (WebQSP) (Yih et al., 2016) and Complex WebQuestions (CWQ) (Talmor & Berant, 2018). These datasets consist of natural language questions designed to be answered using a KG. The questions are open-domain and require retrieval of structured information from Freebase (Bollacker et al., 2008), a large-scale KG with 88 million entities and 126 million triples. Each question in these datasets is annotated with a topic entity and an answer entity.
Topic Entity: This is the primary entity mentioned in the question, serving as the starting point for retrieving triples from the KG. In the running example, “JustinBieber” is the topic entity.
Answer Entity: This is the specific entity within the KG that answers the given question. In our running example, the answer entity would be “JaxonBieber”, representing Justin Bieber’s brother. Note that we might have multiple answer entities for the same question.
2.3 Deletion Strategies
To evaluate the robustness of three popular KG-RAG methods (TOG (Sun et al., 2024), ROG (Luo et al., 2024) and G-Retrieve (He et al., 2024)) under KG incompleteness, we simulate missing knowledge by systematically removing triples from the KG of WebQSP and CWQ. For ToG we choose LLama2-70B-Chat as backbone model. We employ two different deletion strategies: random triple deletion and reasoning path disruption, each designed to model different levels of KG incompleteness.
Random Triple Deletion.
We randomly remove a specified percentage of triples from the KG. This approach provides a baseline for understanding the general impact of missing knowledge. By varying the deletion percentage, we analyze how KG-based RAG performance degrades as more information is lost.
Reasoning Path Disruption.
To simulate more severe knowledge incompleteness, this strategy selectively disrupts reasoning paths that support LLMs’ inference. For each question, we identify the shortest reasoning paths between the topic and answer entities using the breadth-first search algorithm (Moore, 1959). We then randomly select one such path and remove a randomly chosen triple from it. This deliberately breaks a critical step in the reasoning chain, emulating real-world scenarios where essential intermediate knowledge may be missing.
2.4 Evaluation Metrics
The performance of KG-RAG is evaluated by Accuracy and Hits. Let be the set of test questions. For each question , let be the set of tokens generated from the prompt and be the corresponding set of ground truth answer tokens. We define:
3 Experimental Results
Table 1 presents the performance of three KG-RAG methods under varying levels of KG incompleteness. We evaluate models in multiple settings: a fully intact KG (no deletion), random triple deletion at 5%, 10%, and 20% levels, reasoning path disruption (disabling a key step in the reasoning chain), and a baseline condition where retrieval from the KG is entirely disabled. Our findings highlight three key observations.
First, using KG as retrieval source significantly improves performance compared to the LLM without retrieval. For example, RoG’s accuracy falls from 76.75% to 50.46% (-34.25%) in WebQSP. However, all evaluated KG-RAG methods exhibit high sensitivity to missing knowledge in the KG. Notably, even under random deletion, where test question-specific triples may remain unaffected, accuracy consistently declines as the proportion of deleted triples increases across both datasets. The relative accuracy drop is up to 5.99% in WebQSP and 7.41% in CWQ under 20% random triple deletion.
Second, disrupting a single reasoning path leads to a substantial performance drop, despite that multiple reasoning paths often exist for a given question. For instance, in WebQSP, RoG’s accuracy drops from 76.75% to 65.43% (-14.70%), and its hits declines from 86.6% to 78.12% (-9.80%). This suggests that current KG-RAG methods often rely on specific paths for reasoning, rather than effectively utilizing alternative reasoning pathways when available. Models required fine-tuning on the KG such as RoG may be particularly dependent on a small set of learned retrieval patterns, making them vulnerable to missing knowledge in those specific paths.
Finally, our results indicate that KGs remain valuable retrieval sources even when incomplete. Although performance degrades under KG incompleteness, it still surpasses the baseline where no KG retrieval is used. This suggests that addressing KG incompleteness more effectively—rather than discarding KGs entirely—should be the focus when designing more resilient KG-RAG methods.
| Experimental Setting | WebQSP | CWQ | ||
| Accuracy (%) | Hits (%) | Accuracy (%) | Hits (%) | |
| RoG - w/o deletion | 76.75 | 86.60 | 57.49 | 61.79 |
| RoG - 5% random deletion | 75.55 (-1.56%) | 85.99 (-0.71%) | 57.33 (-0.28%) | 61.54 (-0.40%) |
| RoG - 10% random deletion | 74.66 (-2.72%) | 85.68 (-1.06%) | 55.88 (-2.80%) | 60.43 (-2.20%) |
| RoG - 20% random deletion | 72.15 (-5.99%) | 84.15 (-2.83%) | 53.23 (-7.41%) | 58.12 (-5.94%) |
| RoG - reasoning path disruption | 65.43 (-14.70%) | 78.12 (-9.80%) | 52.95 (-7.90%) | 57.4 (-7.10%) |
| RoG - w/o retrieval | 50.46 (-34.25%) | 65.05 (-24.90%) | 35.40 (-38.42%) | 40.10 (-35.11%) |
| ToG - w/o deletion | 44.19 | 67.40 | 48.94 | 51.80 |
| ToG - 5% random deletion | 43.74 (-1.02%) | 67.17 (-0.33%) | 48.78 (-0.32%) | 51.34 (-0.89%) |
| ToG - 10% random deletion | 43.09 (-1.64%) | 66.81 (-0.88%) | 47.94 (-2.05%) | 50.44 (-2.63%) |
| ToG - 20% random deletion | 41.61 (-3.42%) | 66.18 (-1.81%) | 46.14 (-5.73%) | 49.18 (-5.05%) |
| ToG - reasoning path disruption | 38.25 (-13.44%) | 62.70 (-6.97%) | 44.56 (-8.95%) | 47.31 (-8.67%) |
| ToG - w/o retrieval | 34.63 (-21.60%) | 57.24 (-15.08%) | 34.39 (-29.73%) | 37.40 (-27.80%) |
| G-Retriever - w/o deletion | 53.43 | 71.49 | 35.39 | 41.51 |
| G-Retriever - 5% random deletion | 52.33 (-2.06%) | 70.91 (-0.81%) | 34.63 (-2.15%) | 40.55 (-2.31%) |
| G-Retriever - 10% random deletion | 51.87 (-2.92%) | 70.76 (-1.02%) | 33.98 (-3.98%) | 40.27 (-2.99%) |
| G-Retriever - 20% random deletion | 50.88 (-4.78%) | 69.70 (-2.51%) | 33.78 (-4.55%) | 39.98 (-3.69%) |
| G-Retriever - reasoning path disruption | 49.36 (-7.61%) | 69.41 (-2.91%) | 33.56 (-5.17%) | 39.73 (-4.29%) |
| G-Retriever - w/o retrieval | 29.72 (-44.37%) | 47.17 (-34.01%) | 14.86 (-58.01%) | 17.55 (-57.72%) |
4 Conclusion and Future Work
This work evaluated three KG-RAG methods under knowledge incompleteness, demonstrating their sensitivity to missing information. Even random triple deletions reduced performance, while reasoning path disruption led to substantial performance drops. Despite this, KGs remained valuable retrieval sources, consistently outperforming retrieval-free baselines.
Beyond incompleteness, real-world KGs also contain noise, such as incorrect triples generated during automatic KG construction process or errors in the original knowledge source (Heindorf et al., 2016; Paulheim, 2016). Additionally, the lack of a standardized evaluation protocol for KG-RAG methods and the inherent factual instability of LLMs (Potyka et al., 2024; He et al., 2025) can lead to inconsistent assessments across studies.
Future work should focus on developing more robust KG-RAG methods that can handle both missing and noisy knowledge, ensuring reliable performance in real-world scenarios. This includes uncertainty-aware retrieval, noise-resistant reasoning mechanisms, and hybrid approaches that integrate structured and unstructured sources. Additionally, creating comprehensive benchmarks and standardized evaluation protocols will enable systematic assessment and foster progress toward more resilient KG-RAG systems.
5 Acknowledgements
The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Yuqicheng Zhu. The work was partially supported by EU Projects Graph Massivizer (GA 101093202), enRichMyData (GA 101070284) and SMARTY (GA 101140087).